# It’s Not Easy Being Green: On the Energy Efficiency of Programming Languages

### v2

@@ Abstract.
Programming Languages, Performance, Sustainability
ccs: Software and its engineering General programming languagesccs: Software and its engineering Software performanceccs: Social and professional topics Sustainability
1. Introduction
The acceleration of climate change due to use of fossil fuels has driven an increased focus on efforts to decrease both the energy consumption and carbon footprint of computer systems (Hanafy et al., 2024b, a; Radovanovic et al., 2021). In 2018, an estimated 1% of total global energy consumption was attributed to datacenters alone (Masanet et al., 2020). Modern machine learning workloads—especially training—can generate hundreds of tons of CO2 emissions (Strubell et al., 2019; Patterson et al., 2021). For example, Meta reports that training the Llama2 large language model generated an estimated 539 tons of CO2 emissions (GenAI, 2023).
According to an influential line of work, one potential way to reduce energy consumption is to choose a different programming language. This work analyzes a wide selection of programming languages and workloads, and concludes that different programming languages consume widely varying amounts of energy (Couto et al., 2017; Pereira et al., 2017, 2021). The centerpiece of these papers is a ranking of programming languages by energy efficiency (reproduced in part in Table 1).
@@ 1. Introduction
Figure 1. Prior work presented in industry talks to advocate for Rust over other languages due to reduced energy consumption (§2.1).
[…]
Screenshots of video recordings from industry talks showcasing results from prior work.
These studies have received wide attention, both in academic and industrial circles. They have been collectively cited nearly 700 times per Google Scholar (as of March 2025). The results—especially the ranking of programming languages—have had an unusually visible impact in industry, and are routinely quoted on social networks and in blog posts. As Figure 1 illustrates, the rankings have been cited by executives and engineers from Amazon (Miller and Lerche, 2022, 2021; Miller, 2022; Vogels, 2023), Intel (Mulhall and Collins, 2022), SAP (Muthaiah, 2022), and other companies (Suárez et al., 2021) to argue for business decisions and to advocate for a shift in programming languages with an eye towards sustainability. Often, these rankings have been harnessed to support the adoption of Rust, which ranks as one of the most energy-efficient languages while providing safety guarantees that languages like C and C++ lack.
Despite the fact that these studies are statistical and only establish associations, they have nonetheless been broadly interpreted as establishing a causal relationship, that the choice of programming language has a direct effect on a system’s energy consumption. This misinterpretation stems in part from the work’s presentation, not only in ranking of languages by efficiency, but also from the specific claim that “it is almost always possible to choose the best language” when considering execution time and energy consumption (Pereira et al., 2021, §3.3).
Figure 2. The causal model (represented as a causal diagram) of the relationship between programming language and energy consumption presented in this paper (§4). Gray arrows represent comparatively weaker relationships, as Section 4 details: programming language implementation has only a minor effect on parallelism, and memory activity plays a minimal, less controllable role in energy consumption compared to CPU activity.
[…]
A graph with eight nodes: “Programming Language”, “Programming Language Implementation”, “Application Implementation”, “Number of Active Cores”, “Memory Activity”, “Power”, “Time”, and “Energy” There are normal edges: from “Programming Language” to “Programming Language Implementation” and “Application Implementation”; from “Programming Language Implementation” to “Memory Activity” and “Time”; from “Application Implementation” to “Number of active cores”, “Memory Activity” and “Time”; from “Number of Active Cores” to “Power”; from “Power” to “Energy”; from “Time” to “Energy”. There are gray edges: from “Programming Language Implementation” to “Number of Active Cores”; from “Memory Activity” to “Power”.
@@
The above analysis and approach suffer from numerous other methodological flaws (§2.2, §3, §4.2, §4.5). However, the primary focus of this paper is carefully addressing the question: does the choice of programming language affect energy consumption? We embrace this question by developing a rich causal model of the relationship between programming language and energy consumption. Figure 2 presents a causal diagram representing of our final model. Causal diagrams illustrate how different factors influence each other, with arrows showing the direction of these influences (Pearl, 2009). We build this model incrementally, incorporating the various factors at play and their relationships, and provide quantitative and/or qualitative evidence supporting each step.
@@
Our model specifically identifies the number of active cores and memory activity as the key contributors to power draw and consequently energy consumption. It highlights the importance of both programming language and application implementations: language implementation decisions can add execution time overheads including garbage collection or just-in-time warmup, while application implementations dictate the level of parallelism and overall performance of the program. These varying language and application implementation characteristics are the primary drivers of the anomalies we identify in prior work.
@@
Our experimental approach incorporates several methodological improvements, correcting technical errors in prior work that led to negative or otherwise incorrect energy readings. We present an improved methodology based on hardware performance counter data that accurately quantifies average core usage and memory activity. When controlling for confounds like varying CPU utilization and normalizing by time, we conclude that energy consumed is directly proportional to execution time, and independent of the choice of programming language. These results suggest that, to minimize energy consumption, programmers should focus primarily on optimizing performance.
@@
The remainder of this paper is organized as follows: Section 2 introduces and presents a critical analysis of prior work. Section 3 describes the improved methodology we adopt in our experiments, detailing the measurement tool and the benchmarks used. Section 4 incrementally builds the causal model of the relationship between programming language and energy consumption, explaining and correcting anomalies found in prior work at each step. Section 5 concludes with a discussion of other related work.
2. Prior Work
2.1. Overview
@@
The line of work that is the point of departure for this paper explores the relationship between the choice of programming language and energy efficiency (Couto et al., 2017; Pereira et al., 2017, 2021). These papers rank programming languages based on energy consumption, execution time, and memory usage, and attempt to find associations between these metrics. In the remainder of this paper, we refer to these papers collectively as “Pereira et al.”.
@@
Table 1. Partial results from Pereira et al. (§2.1), showing average execution time and energy consumption for the languages discussed in this paper, normalized to C. Boldface denotes anomalous results that Sections 2.2, 4.2, and 4.5 address.
@@
In essence, these studies compare the execution time, energy consumption and memory usage of benchmark implementations in 27 various programming languages. Table 1 presents some of the main results from Pereira et al., limited to the languages discussed in this paper, with anomalous results highlighted in boldface (Section 2.2 discusses this selection in detail). Their experiments leverage Intel’s Running Average Power Limit (RAPL) interface to measure energy consumption, and GNU time or Python’s memory_profiler to measure peak or “total” memory usage, respectively.
The programs used for comparison are from the Computer Language Benchmark Game (Gouy, 2024) (CLBG), a corpus of small benchmark implementations in various languages (22 to 287 lines of code). The most recent paper in the series (Pereira et al., 2021) also incorporates 9 benchmarks from Rosetta Code, a repository of even smaller and simpler code snippets such as Fibonacci, Ackermann, or Sieve of Eratosthenes (typically under 50 lines of code). Because Rosetta Code benchmarks are strictly smaller in scope and size than the CLBG benchmarks, we exclude them from consideration in this paper.
These studies first identify a strong relationship between energy and time. Of course, energy is a linear function of time, hence a strong correlation is expected. They next investigate whether a fast language is always more energy efficient, and claim that this is not the case. They report no significant correlation between peak memory usage and energy consumption, but a strong correlation when considering total memory usage. Finally, they conclude by presenting a ranking of programming languages by energy efficiency, execution time, and memory usage, which programmers can use to select a language for their project.
2.2. Critique
@@ 2.2.1. Programming Language versus Implementation
Programming languages define the syntax and semantics, but it is their implementations that primarily influence performance. While some languages have only a single, widely-used implementation such as Rust or Go, others have multiple implementations, each with their own performance characteristics. For instance, Pereira et al. treat Ruby and JRuby as different languages, while they are in fact two separate implementations of the same Ruby language. The papers uses different benchmark implementations to compare these two Ruby implementations, confounding their comparison.
2.2.2. Selection of Benchmark Implementations
@@
While benchmark implementations are claimed to employ the “exact same algorithm” across languages (Pereira et al., 2021, §1), this is in fact not the case. Benchmark implementations notably have highly varied levels of parallelism and CPU usage, varying degrees of use of third-party libraries, and non-uniform use of vector instructions. These important differences are not properties of the languages themselves, and their effect on performance and energy consumption must be accounted for.
2.2.3. Apparent Anomalies
@@
Some results reported by Pereira et al. are counter-intuitive, and are presented without investigation or explanation. C++ is reported as being 34% less energy efficient and 56% slower than C. Since C++ is approximately a superset of C, and both share the same compiler, optimizations, and code generation backend, we would expect identical energy consumption and execution time. Similarly, TypeScript is reported as being 4.8×4.8$\times$4.8 × less energy efficient and 7.1×7.1$\times$7.1 × slower than JavaScript. TypeScript is a strict superset of JavaScript: any valid JavaScript program is a valid TypeScript program. The same Node.js runtime system is used for both languages. The compilation process for TypeScript may insert or modify code to support older JavaScript standards, but we do not expect this to result in any significant performance overhead. Java, C#, and Go are reported as 1.89×1.89$\times$1.89 ×, 3.14×3.14$\times$3.14 ×, and 2.83×2.83$\times$2.83 × slower than C, respectively. These numbers are unexpectedly high as these implementations are known to be highly optimized (Lion et al., 2022). We expect low overhead from garbage collection costs, plus initial just-in-time compilation overhead for Java and C#. Lua and TypeScript both stand out as exhibiting significantly lower normalized energy consumption than normalized execution time; this is explained by the fact that nearly all of the Lua and TypeScript benchmark implementations are sequential, while the implementations of these benchmarks in other languages are parallelized. Section 4.4.1 characterizes the impact of the number of active cores on energy consumption.
2.2.4. Memory Metrics
@@
Peak memory usage and “total” memory usage are used to estimate memory activity. The tools used to gather both metrics are both based on resident set size (RSS), which is a poor proxy for memory usage and activity (Berger et al., 2023). RSS includes memory that may not be actively used by the program, and crucially does not take cache activity into account. A well-optimized program can have a large RSS but excellent cache locality: if the cache can fulfill most memory requests, memory activity will be low. Benchmark implementation specifics heavily influence peak memory usage, which depends primarily on the choice of data structures used throughout the program. These differences should not be attributed to the languages themselves.
@@
Total memory usage is measured by summing each measurement of RSS taken using Python’s memory_profiler, which is sampled at a frequency of ten times per second. This metric is directly proportional to and thus a proxy for the benchmark’s execution time, which is itself directly proportional to energy consumption.
2.2.5. Language Implementation Specifics
@@
Most language implementations have an initial cost to import and set up necessary in-memory data structures. Language implementations using a just-in-time compiler also require an initial warmup period (Traini et al., 2023; Barrett et al., 2017). These start-up costs are amplified by the benchmarks’ short execution times, which in some cases run for less than a fraction of a second. For example, measuring only the first iteration of a short-lived Java benchmark may not be indicative of Java’s overall performance. Section 4.2.1 shows that this first iteration can be up to 3×3$\times$3 × slower than subsequent ones. Garbage collection also can have a significant impact on both execution time and memory usage (Hertz and Berger, 2005), and can be fine-tuned to obtain better performance.
3. Methodology
Figure 3. Overview of the energy measurement tool we develop for this paper (§3.2). The tool reads performance counters at the beginning and end of each benchmark run, and samples energy consumption readings at regular intervals to avoid overflow. Once the program exits, it saves performance counter values and final energy consumption readings to a file. Periodic sampling and minor post-processing are necessary to ensure correct energy readings, as detailed in Section 3.2.1.
[…]
A diagram depicting step-by-step execution: 1. “Reset performance counters”, 2. “Start program” and “Continuously read RAPL MSRs”, 3. “Program exit”, signaling to stop reading MSRs, 4. “Read performance counters”, and 5. “Synthesize and save measurements”.
@@ 3. Methodology
Table 2. Versions of compilers and runtime systems used (§3.1). All are most recent as of September 2024.
Language Implementation Language(s) Version
@@
3.1. Setup
We conduct experiments in this paper using a server equipped with two Intel Xeon Gold 6430 processors totaling 128 logical cores and 128GB of memory, running Linux version 6.8.0-45-generic. Table 2 details specific compiler and runtime system versions. Experiments in both this paper and Pereira et al.’s use a server architecture.
Benchmark runs are isolated in Docker containers to ensure reproducibility. Docker introduces negligible performance overhead (Felter et al., 2015), which we confirm for the experiments detailed in this paper by obtaining equivalent results with and without Docker. Docker also provides easy access to control groups, making it possible to place restrictions on CPU usage or to pin programs to execute on specific cores. Section 4.3 details how we eliminate the effect of varying concurrency in different benchmark implementations by limiting benchmarks to execute on a single core.
3.2. Measurement tool
Figure 3 provides an overview of the general structure of our energy measurement tool. Our tool takes energy samples once per second so as to minimize overhead.
3.2.1. RAPL
Since the Sandy Bridge microarchitecture (2011), most Intel processors provide a power management interface called Running Average Power Limit (RAPL). RAPL allows measurement of the energy consumption of various parts of the system (Khan et al., 2018). The interface is exposed via Model Specific Registers (MSRs), of which the following measure energy consumption:
  1. (1)
@@ 3.2.1. RAPL
An inspection of Pereira et al.’s data files (Couto and Pereira, 2017) reveals the presence of multiple negative energy readings. Of course, negative energy consumption is not physically possible. We expect that these values were removed by the outlier removal process described in their work and did not affect final results. To ensure correct energy readings, we developed a new measurement tool that corrects the following errors in the tool used by previous work:
  1. (1)
    The previous tool incorrectly includes the upper 32 bits of the energy status counters in their accounting. These upper bits are reserved and must be discarded; our tool tracks only the bottom 32 bits, which contain the actual count of consumed energy.
  2. (2)
    The previous tool only reads a start and end value of the energy status counter, but that counter overflows after “around 60 seconds when power consumption is high” (Intel, 2024). To avoid overflow for long-running computations, our tool uses a separate thread that periodically (at 1Hz) reads the contents of this register.
  3. (3)
    The previous tool incorrectly subtracts counter readings after scaling and conversion to floating point numbers. This conversion will result in negative or otherwise incorrect results when overflow occurs. Our tool avoids this problem by converting results to floating point only after subtracting consecutive readings.
3.2.2. Performance counters

### v3

@@ Abstract
11footnotetext: Work done at the University of Massachusetts Amherst.
I Introduction
The acceleration of climate change due to use of fossil fuels has driven an increased focus on efforts to decrease both the energy consumption and carbon footprint of computer systems [1, 2, 3]. In 2018, an estimated 1% of total global energy consumption was attributed to datacenters alone [4]. Modern machine learning workloads—especially training—can generate hundreds of tons of CO2 emissions [5, 6]. For example, Meta reports that training the Llama2 large language model generated an estimated 539 tons of CO2 emissions [7].
Programmers care about energy when building applications, but often lack tools to effectively address this concern [8]. According to an influential line of work, one potential way to reduce energy consumption is to choose a different programming language. This work analyzes a wide selection of programming languages and workloads, and concludes that different programming languages consume widely varying amounts of energy [9, 10, 11]. The centerpiece of these papers is a ranking of programming languages by energy efficiency (reproduced in part in Table I).
These studies have received wide attention, both in academic and industrial circles. They have been collectively cited over 730 times per Google Scholar (as of May 2025). The results—especially the ranking of programming languages—have had an unusually visible impact in industry, and are routinely quoted on social networks and in blog posts. As Figure 2 illustrates, the rankings have been cited by executives and engineers from Amazon [12, 13, 14, 15], Intel [16], SAP [17], and other companies [18] to argue for business decisions and to advocate for a shift in programming languages with an eye towards sustainability. Often, these rankings have been harnessed to support the adoption of Rust, which ranks as one of the most energy-efficient languages while providing safety guarantees that languages like C and C++ lack.
Despite the fact that these studies are statistical and only establish associations, they have nonetheless been broadly interpreted as establishing a causal relationship, that the choice of programming language has a direct effect on a system’s energy consumption. This misinterpretation stems in part from the work’s presentation, not only in ranking of languages by efficiency, but also from the specific claim that “it is almost always possible to choose the best language” when considering execution time and energy consumption [11, §3.3].
ProgrammingLanguageProgrammingLanguageImplementationApplicationImplementationNumber ofActive CoresMemoryActivityPowerTimeEnergy
Figure 1: The causal model (represented as a causal diagram) of the relationship between programming language and energy consumption presented in this paper (§IV). Gray arrows represent comparatively weaker relationships, as Section IV details: programming language implementation has only a minor effect on parallelism, and memory activity plays a minimal, less controllable role in energy consumption compared to CPU activity.
@@ I Introduction
Figure 2: Industry keynotes advocating for Rust over other languages due to reduced energy consumption, based on the results of Pereira et al. [9, 10, 11]II-A).
@@
The above analysis and approach suffer from numerous other methodological flaws (§II-B, §III, §IV-B, §IV-E). However, the primary focus of this paper is carefully addressing the question: does the choice of programming language affect energy consumption? We embrace this question by developing a rich causal model of the relationship between programming language and energy consumption. Figure 1 presents a causal diagram representing our final model. Causal diagrams illustrate how different factors influence each other, with arrows showing the direction of these influences [19]. We build this model incrementally, incorporating the various factors at play and their relationships, and provide quantitative and/or qualitative evidence supporting each step.
@@
Our model specifically identifies the number of active cores and memory activity as the key contributors to power draw and consequently energy consumption. It highlights the importance of both programming language and application implementations: language implementation decisions can add execution time overheads including garbage collection or just-in-time warmup, while application implementations dictate the level of parallelism and overall performance of the program. These varying implementation characteristics are the primary drivers of the anomalies we identify in prior work.
@@
Our experimental approach incorporates several methodological improvements, correcting technical errors in prior work that led to negative or otherwise incorrect energy readings. We present an improved methodology based on hardware performance counter data that accurately quantifies average core usage and memory activity. When controlling for confounds like varying CPU utilization, we conclude that energy consumed is directly proportional to execution time, and independent of the choice of programming language. These results simplify the task of programmers who are seeking to achieve higher energy efficiency by focusing their efforts on minimizing execution time.
@@
The remainder of this paper is organized as follows: Section II introduces and presents a critical analysis of prior work. Section III describes the improved methodology we adopt in our experiments, detailing the measurement tool and the benchmarks used. Section IV incrementally builds the causal model of the relationship between programming language and energy consumption, explaining and correcting anomalies found in prior work at each step. Section VI concludes with a discussion of other related work.
II Prior Work
II-A Overview
@@
The line of work that is the point of departure for this paper explores the relationship between the choice of programming language and energy efficiency [9, 10, 11]. These papers rank programming languages based on energy consumption, execution time, and memory usage, and attempt to find associations between these metrics. In the remainder of this paper, we refer to these papers collectively as “Pereira et al.”.
@@
TABLE I: Partial results from Pereira et al. (§II-A), showing average execution time and energy consumption for the languages discussed in this paper, normalized to C. Boldface denotes anomalous results that Sections II-B, IV-B, and IV-E address.
@@
In essence, these studies compare the execution time, energy consumption and memory usage of benchmark implementations in 27 various programming languages. Table I presents some of the main results from Pereira et al., limited to the languages discussed in this paper, with anomalous results highlighted in boldface (Section II-B discusses this selection in detail). Their experiments leverage Intel’s Running Average Power Limit (RAPL) interface to measure energy consumption, and GNU time or Python’s memory_profiler to measure peak or “total” memory usage, respectively.
The programs used for comparison are from the Computer Language Benchmark Game [20] (CLBG), a corpus of small benchmark implementations in various languages (22 to 287 lines of code). The most recent paper in the series [11] also incorporates 9 benchmarks from Rosetta Code, a repository of even smaller and simpler code snippets such as Fibonacci, Ackermann, or Sieve of Eratosthenes (typically under 50 lines of code). Because Rosetta Code benchmarks are strictly smaller in scope and size than the CLBG benchmarks, we exclude them from consideration in this paper.
These studies first identify a strong relationship between energy and time. By definition, energy is a linear function of time (energy = power $\times$ time), hence a strong correlation is to be expected. They next investigate whether a fast language is always more energy efficient, and claim that this is not the case. They report no significant correlation between peak memory usage and energy consumption, but a strong correlation when considering total memory usage. Finally, they conclude by presenting a ranking of programming languages by energy efficiency, execution time, and memory usage, which programmers can use to select a language for their project.
II-B Critique
@@
Programming Language versus Implementation: Programming languages define the syntax and semantics, but it is their implementations that primarily influence performance. While some languages have only a single, widely-used implementation such as Rust or Go, others have multiple implementations, each with their own performance characteristics. For instance, Pereira et al. treat Ruby and JRuby as different languages, while they are in fact two separate implementations of the same Ruby language. The papers uses different benchmark implementations to compare these two Ruby implementations, confounding their comparison.
@@
Widely Varying Benchmark Implementations: While benchmark implementations are claimed to employ the “exact same algorithm” across languages [11, §1], this is in fact not the case. Benchmark implementations notably have highly varied levels of parallelism and CPU usage, varying degrees of use of third-party libraries, and non-uniform use of vector instructions. These important differences are not properties of the languages themselves, and their effect on performance and energy consumption must be accounted for.
@@
Apparent Anomalies in Results: Some results reported by Pereira et al. are counter-intuitive, and are presented without investigation or explanation. C++ is reported as being 34% less energy efficient and 56% slower than C. Since C++ is approximately a superset of C, and both share the same compiler, optimizations, and code generation backend, we would expect identical energy consumption and execution time. Similarly, TypeScript is reported as being 4.8×4.8$\times$4.8 × less energy efficient and 7.1×7.1$\times$7.1 × slower than JavaScript. TypeScript is a strict superset of JavaScript: any valid JavaScript program is a valid TypeScript program. The same Node.js runtime system is used for both languages. The compilation process for TypeScript may insert or modify code to support older JavaScript standards, but we do not expect this to result in any significant performance overhead. Java, C#, and Go are reported as 1.89×1.89$\times$1.89 ×, 3.14×3.14$\times$3.14 ×, and 2.83×2.83$\times$2.83 × slower than C, respectively. These numbers are unexpectedly high as these implementations are known to be highly optimized [21]. We expect low overhead from garbage collection costs, plus initial just-in-time compilation overhead for Java and C#. Lua and TypeScript both stand out as exhibiting significantly lower normalized energy consumption than normalized execution time; this is explained by the fact that nearly all of the Lua and TypeScript benchmark implementations are sequential, while the implementations of these benchmarks in other languages are parallelized. Section IV-D1 characterizes the impact of the number of active cores on energy consumption.
@@
Inappropriate Memory Metrics: Peak memory usage and “total” memory usage are used to estimate memory activity. The tools used to gather both metrics are both based on resident set size (RSS), which is a poor proxy for memory usage and activity [22]. RSS includes memory that may not be actively used by the program, and crucially does not take cache activity into account. A well-optimized program can have a large RSS but excellent cache locality: if the cache can fulfill most memory requests, memory activity will be low. Benchmark implementation specifics heavily influence peak memory usage, which depends primarily on the choice of data structures used throughout the program. These differences should not be attributed to the languages themselves.
@@
Total memory usage is measured with Python’s memory_profiler, which samples the RSS at a frequency of ten times per second. This metric is directly proportional to execution time, which is itself directly proportional to energy consumption.
@@
Failure to Control for Language Implementation Specifics: Most language implementations have an initial cost to import and set up necessary in-memory data structures. Language implementations using a just-in-time compiler also require an initial warmup period [23, 24]. These start-up costs are amplified by the benchmarks’ short execution times, which in some cases run for less than a fraction of a second. For example, measuring only the first iteration of a short-lived Java benchmark may not be indicative of Java’s overall performance. Section IV-B1 shows that this first iteration can be up to 3×3$\times$3 × slower than subsequent ones. Garbage collection also can have a significant impact on both execution time and memory usage [25], and can be fine-tuned to obtain better performance.
III Methodology
@@
TABLE II: Versions of compilers and runtime systems used (§III-A). All are most recent as of September 2024.
Implementation Language(s) Version
@@
III-A Setup
We conduct experiments in this paper using a server equipped with two Intel Xeon Gold 6430 processors totaling 128 logical cores and 128GB of memory, running Linux version 6.8.0-45-generic. Table II details specific compiler and runtime system versions. Experiments in both this paper and Pereira et al.’s use a server architecture.
Benchmark runs are isolated in Docker containers to ensure reproducibility. Docker introduces negligible performance overhead [26], which we confirm for the experiments detailed in this paper by obtaining equivalent results with and without Docker. Docker also provides easy access to control groups, making it possible to place restrictions on CPU usage or to pin programs to execute on specific cores. Section IV-C details how we eliminate the effect of varying concurrency in different benchmark implementations by limiting benchmarks to execute on a single core.
III-B Measurement Tool
We develop a new measurement tool for this study, correcting errors found in Pereira et al.’s implementation [27], and adding performance counter readings capturing the number of cores used and memory activity. The tool reads performance counters at the beginning and end of each benchmark run, and samples energy consumption readings once per second. Periodic sampling is necessary to ensure correct energy readings, as the next section explains.
III-B1 RAPL
Starting with the Sandy Bridge microarchitecture (2011), most Intel processors provide a power management interface called Running Average Power Limit (RAPL). RAPL allows measurement of the energy consumption of various parts of the system [28]. The interface is exposed via Model Specific Registers (MSRs), of which the following measure energy consumption:
@@ III-B1 RAPL
An inspection of Pereira et al.’s data files [27] reveals the presence of multiple negative energy readings. Of course, negative energy consumption is not physically possible. We expect that these values were removed by the outlier removal process described in their work and did not affect final results. To ensure correct energy readings, we rectify the following errors in the tool used by previous work:
  • Erroneous inclusion of upper bits of status counters: The previous tool incorrectly includes the upper 32 bits of the energy status counters in their accounting. These upper bits are reserved and must be discarded; our tool tracks only the bottom 32 bits, which contain the actual count of consumed energy.
  • Failure to prevent overflow of status counters: The previous tool reads only a start and end value of the energy status counter, but that counter overflows after “around 60 seconds when power consumption is high” [29]. To avoid overflow for long-running computations, our tool uses a separate thread that periodically (at 1Hz) reads the contents of this register.
  • Erroneous floating point arithmetic: Finally, the previous tool incorrectly subtracts counter readings after scaling and conversion to floating point. This conversion will result in negative or otherwise incorrect results when overflow occurs. Our tool avoids this problem by converting results to floating point only after subtracting consecutive readings.
III-B2 Performance Counters